# A Survey of Reinforcement Learning For Economics

### 17 Mar 2026 v3

@@ 6.7 Simulation Study: Bellman Error and Value Error in Offline Policy Evaluation
Two algorithms are trained on $\mathcal{D}$.11911911950,000 gradient steps per seed, 3 seeds; two-layer MLP with 64 hidden units, Adam at $5\times 10^{-4}$; target network updated every 500 steps. BRM minimizes $\mathbb{E}_{\mathcal{D}}[(Q(s,a)-(r+\gamma\max_{a^{\prime}}Q(s^{\prime},a^{\prime})))^{2}]$ where both $Q(s,a)$ and $Q(s^{\prime},a^{\prime})$ use the current network; gradients flow through both sides simultaneously. The key consequence is that the network can zero the residual at an observed pair $(s,a)$ by co-moving $Q(s,a)$ and $Q(s^{\prime},a^{\prime})$ together, rather than moving either toward $Q^{*}$. This is the opposite of supervised learning, where labels are fixed external targets that do not move with the model weights. FQE prevents this by using a frozen target network for $Q(s^{\prime},a^{\prime})$; the network must reduce $Q(s,a)$ toward a target that does not respond to its own gradient steps. Every 500 steps we record the Bellman error on $\mathcal{D}$ (current network on both sides) and the value error $\frac{1}{|\mathcal{F}|}\sum_{(s,a)\in\mathcal{F}}|Q_{\theta}(s,a)-Q^{*}(s,a)|$ over all 4,795 feasible pairs. As a supervised baseline, OLS regression of $\log c$ on $\log k$ and $\log z$ is estimated on expanding windows.120120120Noise $\sigma=0.30$; windows from $n=10$ to 2,0002{,}000; held-out test set of 500 transitions.
@@ 7.1 Ride-Hailing Dispatch
Qin et al. (2021) formalized DiDi’s order dispatching as a semi-Markov decision process121121121A semi-Markov decision process generalizes the standard MDP by allowing variable time between decisions. The discount factor $\gamma^{\tau_{k}}$ depends on the actual time elapsed $\tau_{k}$ rather than applying a fixed per-step discount. where each driver is an independent agent. The state of a driver consists of location (discretized into hexagonal zones) and time (bucketed into intervals). The action is the order assigned to the driver, with the option to remain idle. The reward is the trip fare. State transitions are determined by trip destinations, as completing an order transports the driver from origin to destination, changing their spatial state. The stochasticity arises from future demand, which determines the available actions at each state. The per-driver value function $V^{\pi}(s)$ represents the expected cumulative fare a driver can earn from state $s$ under dispatching policy $\pi$:
@@
where $\hat{r}_{i}$ is the predicted fare for order $i$, $\hat{\tau}_{i}$ is the estimated trip duration, and $s_{j}^{\prime}$ is the destination state. The value function is learned via temporal-difference methods from historical trip data. DiDi’s Cerebellar Value Network (CVNet; Tang et al., 2019) uses hierarchical coarse-coding122122122Coarse-coding represents a value function as a weighted sum over overlapping rectangular or hexagonal tiles that partition the state space (Sutton and Barto, 2018); each state activates the tiles that contain it, and a hierarchical variant stacks tiles at multiple resolutions to allow both coarse and fine generalization simultaneously. with a multi-resolution hexagonal grid, enabling transfer learning123123123Transfer learning initializes a model for a new task (a new city) with parameters trained on related tasks (existing cities), on the assumption that learned representations of traffic and demand patterns generalize across contexts and require less data to reach good performance in the new setting. across cities and robustness to data sparsity in low-traffic zones.
@@
Han et al. (2022) reported complementary results from Lyft, where the dispatching system optimizes driver assignment and repositioning jointly. Their value decomposition architecture124124124Value decomposition decomposes the platform’s global matching objective into individual driver value functions, each estimable via temporal-difference learning. The global optimum is recovered by optimizing the sum of individual values. assigns credit to individual driver decisions within the global objective. The production system demonstrated improvements in rider wait times and driver utilization, providing a second data point suggesting that similar approaches may transfer across platforms.
@@ 7.2 Hotel Revenue Management
Their system uses a two-step design. The RL agent observes the state $(t,s)$, where $t$ indexes the booking period within a $T=10$-period episode and $s$ is the average revenue per room sold to date, and selects an average discount level $a\in\{10\%,20\%,30\%\}$. A linear program then converts this scalar recommendation into a feasible capacity allocation across rate segments, accounting for each hotel’s channel preferences.125125125The RL component uses a modified on-policy Monte Carlo method with $\varepsilon$-greedy exploration, updating Q-values from realized returns after each completed episode. Training on real operational data rather than simulated environments means the method adapts to demand non-stationarity without requiring an explicit demand model. This decomposition addresses practitioner acceptance, since managers understand a single discount recommendation, and circumvents the need to explicitly model demand arrivals or cancellation rates.
The field experiment randomly assigned five Hanting-brand hotels in Shanghai to the treatment group from ten candidates, with 271 additional Shanghai hotels serving as a donor pool for synthetic control estimation.126126126Synthetic control constructs a weighted combination of untreated hotels whose pre-treatment trend matches each treated hotel, avoiding the selection bias of simple before-after comparisons. Table 10 reports the average treatment effects over the pilot period (March–June 2015).
@@ 7.3 E-Commerce Dynamic Pricing
where $\text{uv}_{i,t}$ is unique visitors in period $t$ and $\tau$ is a reference lag.127127127DRCR normalizes revenue by traffic to remove demand fluctuations unrelated to pricing. The differencing further removes product-specific level effects, yielding a reward signal that is more concave than raw revenue and improves convergence stability.
Two algorithms are deployed: DQN for the discrete action formulation and DDPG for continuous actions. Both are pre-trained from demonstrations using historical specialist pricing actions (DQfD, DDPGfD) to address cold-start.128128128Standard A/B testing is infeasible because Chinese e-commerce regulations prohibit displaying different prices to different customers for the same product simultaneously. Liu et al. (2019) instead evaluate using difference-in-differences against “simi-products” (similar products not subject to algorithmic pricing) as controls.
@@ 7.4 Financial Order Execution
Nevmyvaka et al. (2006) applied tabular Q-learning to execution on real limit order book data from NASDAQ stocks.129129129The dataset covers 500 trading days of millisecond-level limit order book snapshots for AMZN, QCOM, and NVDA. The state space contains approximately 10,000 states; the horizon is $T=60$ seconds with discount $\gamma=1$. The state at time $t$ is $s_{t}=(t,q_{t},\psi_{t},\Delta_{t})$, where $t\in\{1,\ldots,T\}$ is time remaining, $q_{t}\in\{0,1,\ldots,Q\}$ is inventory remaining, $\psi_{t}$ is the discretized bid-ask spread, and $\Delta_{t}$ is the discretized signed volume imbalance.130130130Signed volume imbalance is the difference between buy and sell order volume near the best prices, normalized by total volume; positive imbalance typically predicts short-term price increases. Actions $a_{t}\in\{0,\delta,2\delta,\ldots,q_{t}\}$ specify shares to execute in the current interval. The reward is the negative per-period slippage contribution:
@@
where $p_{0}$ is the mid-quote at order arrival and $p_{t}^{\text{exec}}$ is the average execution price. Total implementation shortfall is $IS=-\sum_{t=1}^{T}r_{t}/Q$.131131131Implementation shortfall measures the total cost of executing a trade relative to the benchmark mid-quote at arrival. It captures market impact, timing cost, and opportunity cost of unexecuted shares. Because the objective is cost minimization, the Q-learning update uses $\min$ over next-period actions:
@@ 7.5 Supply Chain Inventory Management
Gijsbrechts et al. (2022) conducted a systematic evaluation of deep RL against classical base-stock policies across lost sales, dual sourcing, and multi-echelon configurations. The classical benchmark is the echelon base-stock policy, in which each stage $k$ maintains an echelon inventory position and orders to bring this position to a target level $S_{k}$.132132132Formally, the echelon inventory position at stage $k$ equals on-hand inventory at $k$ plus all inventory at downstream stages $1,\ldots,k-1$ plus in-transit inventory, minus backorders. The echelon base-stock policy of Clark and Scarf (1960) is optimal for serial systems with linear costs and backorders. For single-echelon systems with backorders, the optimal base-stock level is given by the newsvendor critical fractile $S^{*}=F_{D}^{-1}(b/(b+h))$. Table 13 summarizes results from their multi-echelon experiments.
@@ 6 $\sim 10^{8}$ Failed
Where classical solutions exist, RL underperforms: the base-stock policy remains highly competitive when properly calibrated, and DRL required millions of training transitions to approach performance levels achievable through closed-form calculation. RL struggles particularly with multi-echelon coupling, where upstream orders affect downstream costs many periods later.133133133Credit assignment refers to the difficulty of determining which past actions caused a delayed reward or cost. In multi-echelon systems, an upstream ordering decision today may not affect customer-facing costs for many periods, making it difficult for RL to learn the causal connection. The value proposition for RL lies in problems where analytical solutions are unavailable: non-stationary demand, complex operational constraints, or cost structures that do not admit tractable decomposition. Even in these settings, successful deployment remains rare and requires extensive simulation infrastructure, domain expertise, and careful calibration against classical baselines.
@@ 7.6 Real-Time Bidding
where $\mathrm{cvr}_{t}\in\{0,1\}$ is the conversion indicator. Budget evolves as $B_{t+1}=B_{t}-\mathbbm{1}\{b_{t}\geq c_{t}\}\cdot c_{t}$, and the episode terminates when $B_{t}\leq 0$ or t=0t=0.134134134The state space is discretized into budget bins and time bins. Wu et al. (2018) augment the state with bid landscape features derived from historical clearing price distributions, giving the agent information about current market competitiveness.
@@ 7.7 Simulation Study: Bus Engine Replacement
I extend the single-engine problem to a fleet of $N$ engines with a capacity constraint limiting replacements per period. The state $s=(m_{1},\ldots,m_{N})$ records discretized mileage for each engine. Actions are subsets of engines to replace, subject to the capacity constraint. The per-period cost is $c(s,a)=\alpha\sum_{i}m_{i}+\beta|a|$, where $\alpha$ is the operating cost per unit mileage and $\beta$ is the replacement cost.135135135The mileage-dependent operating cost $\alpha\sum_{i}m_{i}$ follows Rust’s original specification $c(x,\theta_{1})=\theta_{11}x$, creating a non-trivial threshold replacement policy. The fleet extension uses deterministic mileage increments to isolate the combinatorial scaling challenge that arises from the joint state of multiple engines. Mileage evolves deterministically; replaced engines reset to $m=0$; others increment by one bin.136136136With $M=6$ mileage bins, the state space is $6^{N}$: 1,296 states at $N=4$, 7,776 at $N=5$, 46,656 at $N=6$. Value iteration is feasible for $N\leq 5$.
@@ 8 Structural Estimation with Reinforcement Learning
Several recent papers have used RL training loops,137137137Throughout this chapter, the RL training loop runs entirely inside the econometrician’s computational model; the agent never interacts with real economic agents or markets but serves as a numerical method for solving the Bellman equation within a structural estimation procedure (Section 2). There is no execution phase in the usual sense. namely Q-learning, temporal-difference learning, policy gradient, and actor-critic methods, to solve structural economic models at scales where conventional dynamic programming fails.138138138I exclude papers that use neural network function approximation without an RL training mechanism. Inverse reinforcement learning is treated in the sister survey (Rust and Rawat, 2026). Throughout this chapter I adopt a unified notation. An MDP is a tuple $(\mathcal{S},\mathcal{A},P,r,\gamma)$ where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $P(s^{\prime}|s,a)$ is the transition kernel, $r(s,a)$ is the per-period reward, and $\gamma\in[0,1)$ is the discount factor.139139139Several of the papers reviewed here use $\beta$ for the discount factor, following economics convention. I translate all results to $\gamma$ for consistency with the RL literature and the rest of this survey. A policy $\pi:\mathcal{S}\to\Delta(\mathcal{A})$ maps states to distributions over actions. The value function under $\pi$ is $V^{\pi}(s)=\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\mid s_{0}=s\right]$, and the action-value function is $Q^{\pi}(s,a)=r(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim P(\cdot|s,a)}[V^{\pi}(s^{\prime})]$. The optimal value function satisfies $V^{*}(s)=\max_{a\in\mathcal{A}}Q^{*}(s,a)$.
@@
8.0.1 Adusumilli and Eckardt (2022): TD Learning for CCP Estimation

### 23 Mar 2026 v4

@@ 6.7 Simulation Study: Bellman Error and Value Error in Offline Policy Evaluation
Two algorithms are trained on $\mathcal{D}$.12712712750,000 gradient steps per seed, 3 seeds; two-layer MLP with 64 hidden units, Adam at $5\times 10^{-4}$; target network updated every 500 steps. BRM minimizes $\mathbb{E}_{\mathcal{D}}[(Q(s,a)-(r+\gamma\max_{a^{\prime}}Q(s^{\prime},a^{\prime})))^{2}]$ where both $Q(s,a)$ and $Q(s^{\prime},a^{\prime})$ use the current network; gradients flow through both sides simultaneously. The key consequence is that the network can zero the residual at an observed pair $(s,a)$ by co-moving $Q(s,a)$ and $Q(s^{\prime},a^{\prime})$ together, rather than moving either toward $Q^{*}$. This is the opposite of supervised learning, where labels are fixed external targets that do not move with the model weights. FQE prevents this by using a frozen target network for $Q(s^{\prime},a^{\prime})$; the network must reduce $Q(s,a)$ toward a target that does not respond to its own gradient steps. Every 500 steps we record the Bellman error on $\mathcal{D}$ (current network on both sides) and the value error $\frac{1}{|\mathcal{F}|}\sum_{(s,a)\in\mathcal{F}}|Q_{\theta}(s,a)-Q^{*}(s,a)|$ over all 4,795 feasible pairs. As a supervised baseline, OLS regression of $\log c$ on $\log k$ and $\log z$ is estimated on expanding windows.128128128Noise $\sigma=0.30$; windows from $n=10$ to 2,0002{,}000; held-out test set of 500 transitions.
@@ 7.1 Ride-Hailing Dispatch
Qin et al. (2021) formalized DiDi’s order dispatching as a semi-Markov decision process129129129A semi-Markov decision process generalizes the standard MDP by allowing variable time between decisions. The discount factor $\gamma^{\tau_{k}}$ depends on the actual time elapsed $\tau_{k}$ rather than applying a fixed per-step discount. where each driver is an independent agent. The state of a driver consists of location (discretized into hexagonal zones) and time (bucketed into intervals). The action is the order assigned to the driver, with the option to remain idle. The reward is the trip fare. State transitions are determined by trip destinations, as completing an order transports the driver from origin to destination, changing their spatial state. The stochasticity arises from future demand, which determines the available actions at each state. The per-driver value function $V^{\pi}(s)$ represents the expected cumulative fare a driver can earn from state $s$ under dispatching policy $\pi$:
@@
where $\hat{r}_{i}$ is the predicted fare for order $i$, $\hat{\tau}_{i}$ is the estimated trip duration, and $s_{j}^{\prime}$ is the destination state. The value function is learned via temporal-difference methods from historical trip data. DiDi’s Cerebellar Value Network (CVNet; Tang et al., 2019) uses hierarchical coarse-coding130130130Coarse-coding represents a value function as a weighted sum over overlapping rectangular or hexagonal tiles that partition the state space (Sutton and Barto, 2018); each state activates the tiles that contain it, and a hierarchical variant stacks tiles at multiple resolutions to allow both coarse and fine generalization simultaneously. with a multi-resolution hexagonal grid, enabling transfer learning131131131Transfer learning initializes a model for a new task (a new city) with parameters trained on related tasks (existing cities), on the assumption that learned representations of traffic and demand patterns generalize across contexts and require less data to reach good performance in the new setting. across cities and robustness to data sparsity in low-traffic zones.
@@
Han et al. (2022) reported complementary results from Lyft, where the dispatching system optimizes driver assignment and repositioning jointly. Their value decomposition architecture132132132Value decomposition decomposes the platform’s global matching objective into individual driver value functions, each estimable via temporal-difference learning. The global optimum is recovered by optimizing the sum of individual values. assigns credit to individual driver decisions within the global objective. The production system demonstrated improvements in rider wait times and driver utilization, providing a second data point suggesting that similar approaches may transfer across platforms.
@@ 7.2 Hotel Revenue Management
Their system uses a two-step design. The RL agent observes the state $(t,s)$, where $t$ indexes the booking period within a $T=10$-period episode and $s$ is the average revenue per room sold to date, and selects an average discount level $a\in\{10\%,20\%,30\%\}$. A linear program then converts this scalar recommendation into a feasible capacity allocation across rate segments, accounting for each hotel’s channel preferences.133133133The RL component uses a modified on-policy Monte Carlo method with $\varepsilon$-greedy exploration, updating Q-values from realized returns after each completed episode. Training on real operational data rather than simulated environments means the method adapts to demand non-stationarity without requiring an explicit demand model. This decomposition addresses practitioner acceptance, since managers understand a single discount recommendation, and circumvents the need to explicitly model demand arrivals or cancellation rates.
The field experiment randomly assigned five Hanting-brand hotels in Shanghai to the treatment group from ten candidates, with 271 additional Shanghai hotels serving as a donor pool for synthetic control estimation.134134134Synthetic control constructs a weighted combination of untreated hotels whose pre-treatment trend matches each treated hotel, avoiding the selection bias of simple before-after comparisons. Table 10 reports the average treatment effects over the pilot period (March–June 2015).
@@ 7.3 E-Commerce Dynamic Pricing
where $\text{uv}_{i,t}$ is unique visitors in period $t$ and $\tau$ is a reference lag.135135135DRCR normalizes revenue by traffic to remove demand fluctuations unrelated to pricing. The differencing further removes product-specific level effects, yielding a reward signal that is more concave than raw revenue and improves convergence stability.
Two algorithms are deployed: DQN for the discrete action formulation and DDPG for continuous actions. Both are pre-trained from demonstrations using historical specialist pricing actions (DQfD, DDPGfD) to address cold-start.136136136Pre-populating replay buffers with rollouts from heuristic or scripted policies is a general technique for reducing exploration burden in deep RL. In robotic grasping, Kalashnikov et al. (2018) collected 200,000 scripted grasping attempts at 15–30% success rate, sufficient to bootstrap a vision-based policy to 96% success; see Ibarz et al. (2021) for a survey of these and related warm-start strategies.137137137Standard A/B testing is infeasible because Chinese e-commerce regulations prohibit displaying different prices to different customers for the same product simultaneously. Liu et al. (2019) instead evaluate using difference-in-differences against “simi-products” (similar products not subject to algorithmic pricing) as controls.
@@ 7.4 Financial Order Execution
Nevmyvaka et al. (2006) applied tabular Q-learning to execution on real limit order book data from NASDAQ stocks.138138138The dataset covers 500 trading days of millisecond-level limit order book snapshots for AMZN, QCOM, and NVDA. The state space contains approximately 10,000 states; the horizon is $T=60$ seconds with discount $\gamma=1$. The state at time $t$ is $s_{t}=(t,q_{t},\psi_{t},\Delta_{t})$, where $t\in\{1,\ldots,T\}$ is time remaining, $q_{t}\in\{0,1,\ldots,Q\}$ is inventory remaining, $\psi_{t}$ is the discretized bid-ask spread, and $\Delta_{t}$ is the discretized signed volume imbalance.139139139Signed volume imbalance is the difference between buy and sell order volume near the best prices, normalized by total volume; positive imbalance typically predicts short-term price increases. Actions $a_{t}\in\{0,\delta,2\delta,\ldots,q_{t}\}$ specify shares to execute in the current interval. The reward is the negative per-period slippage contribution:
@@
where $p_{0}$ is the mid-quote at order arrival and $p_{t}^{\text{exec}}$ is the average execution price. Total implementation shortfall is $IS=-\sum_{t=1}^{T}r_{t}/Q$.140140140Implementation shortfall measures the total cost of executing a trade relative to the benchmark mid-quote at arrival. It captures market impact, timing cost, and opportunity cost of unexecuted shares. Because the objective is cost minimization, the Q-learning update uses $\min$ over next-period actions:
@@ 7.5 Supply Chain Inventory Management
Gijsbrechts et al. (2022) conducted a systematic evaluation of deep RL against classical base-stock policies across lost sales, dual sourcing, and multi-echelon configurations. The classical benchmark is the echelon base-stock policy, in which each stage $k$ maintains an echelon inventory position and orders to bring this position to a target level $S_{k}$.141141141Formally, the echelon inventory position at stage $k$ equals on-hand inventory at $k$ plus all inventory at downstream stages $1,\ldots,k-1$ plus in-transit inventory, minus backorders. The echelon base-stock policy of Clark and Scarf (1960) is optimal for serial systems with linear costs and backorders. For single-echelon systems with backorders, the optimal base-stock level is given by the newsvendor critical fractile $S^{*}=F_{D}^{-1}(b/(b+h))$. Table 13 summarizes results from their multi-echelon experiments.
@@ 6 $\sim 10^{8}$ Failed
Where classical solutions exist, RL underperforms: the base-stock policy remains highly competitive when properly calibrated, and DRL required millions of training transitions to approach performance levels achievable through closed-form calculation. RL struggles particularly with multi-echelon coupling, where upstream orders affect downstream costs many periods later.142142142Credit assignment refers to the difficulty of determining which past actions caused a delayed reward or cost. In multi-echelon systems, an upstream ordering decision today may not affect customer-facing costs for many periods, making it difficult for RL to learn the causal connection. The value proposition for RL lies in problems where analytical solutions are unavailable: non-stationary demand, complex operational constraints, or cost structures that do not admit tractable decomposition.143143143Residual RL (Silver et al., 2018b) offers a middle ground: rather than replacing classical policies, learn a correction $\pi_{\mathrm{final}}(s)=\pi_{\mathrm{classical}}(s)+\pi_{\mathrm{learned}}(s)$, preserving the classical solution’s performance while allowing RL to handle constraints or non-stationarity the analytical method cannot. This connects to perturbation methods in economics, where one linearizes around a known solution and solves for deviations. See Ibarz et al. (2021) for a survey of residual and demonstration-augmented RL in robotics. Even in these settings, successful deployment remains rare and requires extensive simulation infrastructure, domain expertise, and careful calibration against classical baselines.
@@ 7.6 Real-Time Bidding
where $\mathrm{cvr}_{t}\in\{0,1\}$ is the conversion indicator. Budget evolves as $B_{t+1}=B_{t}-\mathbbm{1}\{b_{t}\geq c_{t}\}\cdot c_{t}$, and the episode terminates when $B_{t}\leq 0$ or t=0t=0.144144144The state space is discretized into budget bins and time bins. Wu et al. (2018) augment the state with bid landscape features derived from historical clearing price distributions, giving the agent information about current market competitiveness.
@@ 7.7 Simulation Study: Bus Engine Replacement
I extend the single-engine problem to a fleet of $N$ engines with a capacity constraint limiting replacements per period. The state $s=(m_{1},\ldots,m_{N})$ records discretized mileage for each engine. Actions are subsets of engines to replace, subject to the capacity constraint. The per-period cost is $c(s,a)=\alpha\sum_{i}m_{i}+\beta|a|$, where $\alpha$ is the operating cost per unit mileage and $\beta$ is the replacement cost.145145145The mileage-dependent operating cost $\alpha\sum_{i}m_{i}$ follows Rust’s original specification $c(x,\theta_{1})=\theta_{11}x$, creating a non-trivial threshold replacement policy. The fleet extension uses deterministic mileage increments to isolate the combinatorial scaling challenge that arises from the joint state of multiple engines. Mileage evolves deterministically; replaced engines reset to $m=0$; others increment by one bin.146146146With $M=6$ mileage bins, the state space is $6^{N}$: 1,296 states at $N=4$, 7,776 at $N=5$, 46,656 at $N=6$. Value iteration is feasible for $N\leq 5$.
@@ 8 Structural Estimation with Reinforcement Learning
Several recent papers have used RL training loops,147147147Throughout this chapter, the RL training loop runs entirely inside the econometrician’s computational model; the agent never interacts with real economic agents or markets but serves as a numerical method for solving the Bellman equation within a structural estimation procedure (Section 2). There is no execution phase in the usual sense. namely Q-learning, temporal-difference learning, policy gradient, and actor-critic methods, to solve structural economic models at scales where conventional dynamic programming fails.148148148I exclude papers that use neural network function approximation without an RL training mechanism. Inverse reinforcement learning is treated in the sister survey (Rust and Rawat, 2026). Throughout this chapter I adopt a unified notation. An MDP is a tuple $(\mathcal{S},\mathcal{A},P,r,\gamma)$ where $\mathcal{S}$ is the state space, $\mathcal{A}$ is the action space, $P(s^{\prime}|s,a)$ is the transition kernel, $r(s,a)$ is the per-period reward, and $\gamma\in[0,1)$ is the discount factor.149149149Several of the papers reviewed here use $\beta$ for the discount factor, following economics convention. I translate all results to $\gamma$ for consistency with the RL literature and the rest of this survey. A policy $\pi:\mathcal{S}\to\Delta(\mathcal{A})$ maps states to distributions over actions. The value function under $\pi$ is $V^{\pi}(s)=\mathbb{E}_{\pi}\left[\sum_{t=0}^{\infty}\gamma^{t}r(s_{t},a_{t})\mid s_{0}=s\right]$, and the action-value function is $Q^{\pi}(s,a)=r(s,a)+\gamma\mathbb{E}_{s^{\prime}\sim P(\cdot|s,a)}[V^{\pi}(s^{\prime})]$. The optimal value function satisfies $V^{*}(s)=\max_{a\in\mathcal{A}}Q^{*}(s,a)$.
@@
8.1 Single-Agent Structural Estimation
8.1.1 TD Learning for CCP Estimation